### v1

# Contemporaneous and lagged spillovers across crude oil, carbon emission allowance, climate change, and agriculture futures markets: Evidence from the $R^{2}$ decomposed connectedness approach

Abstract
In this paper, we examine the dynamic spillovers among the crude oil, carbon emission allowance, climate change, and agricultural markets. Adopting a novel $R^{2}$ decomposed connectedness approach, our empirical analysis reveals several key findings. The overall TCI dynamics have been mainly dominated by contemporaneous dynamics rather than the lagged dynamics. We also find climate change has significant spillovers to other markets. Moreover, there are heterogeneous spillover effects among agricultural markets. Specially, corn is the biggest risk contributor to this system, while barley is the major risk receiver of shocks.
keywords:
Dynamic connectedness , Crude oil , Carbon emission allowance , Climate change , Agricultural commodities , $R^{2}$ decomposition
journal: Elsevier
1 Introduction
As global warming intensifies and extreme weather becomes more frequent, concerns about potential economic losses and impacts on social development have increased, leading to widespread attention to climate risks (Venturini, 2022; Kemp et al., 2022; Pham and Kamal, 2024; Khalfaoui et al., 2024). Severe physical climate risks affect the cultivation of agricultural products. Additionally, carbon dioxide also exerts a negative impact on agricultural and food markets (Khalfaoui et al., 2024). Climate change influences the extraction and transportation of fossil energy. Volumetric oil production declines with significant growth of greenhouse gas (Masnadi and Brandt, 2017). However, energy production and agricultural activities can lead to environmental pollution, worsening climate conditions and intensifying global warming (Jin et al., 2023).
Since agricultural and energy markets have always been fundamentally connected (Du et al., 2011; Shao et al., 2024), climate change has added new complexity to this interconnection. Moreover, the financialization of commodities has enriched investment instruments but has also heightened the risk of contagion. Owing to the strong interconnections within the financial system, a shock in one sector can affect other sectors. Furthermore, the contagion effect can extend to the entire financial market. Hence, the research of spillover effect and information transmission among climate change and commodity markets continues to expand, which is crucial for assessing contagion risks, price forecasting, and portfolio management. Jin et al. (2023) report the dynamic spillover among geopolitical risk, climate risk, and energy markets from an international perspective. Lin et al. (2024) observe simultaneous opposite correlations between climate change attention and crude oil markets.
In response to the crucial climate issues, a carbon emission allowance trading market has been developed. The market provides economic incentives for companies to adopt cleaner technologies and energy sources to reduce emission costs, helping to mitigate climate change. It should be noted that fossil energy and agricultural markets have important impacts on the carbon market, which will influence carbon demand, supply, and prices (Gong et al., 2021). Chen et al. (2024) provide empirical evidence that the risk is transferred from the international gas and crude oil markets to the Chinese carbon market. Younis et al. (2024) investigates contagion risk among climate change, renewable energy, technological innovation, and G-17 banking stock markets. They find significant risk spillover and connectedness.
Recently, the interdependency among energy and agricultural commodity markets has been widely concerned (Wei et al., 2023). However, existing research has mainly focused on overall spillover effects as well as long-term and short-term frequency spillovers. There is relatively limited empirical analysis that decomposes the spillover effects into contemporaneous and lagged periods (Balli et al., 2023). Furthermore, there is a lack of detailed exploration of the risk spillover among agricultural markets in existing studies. Hence, we examine the dynamic spillover among carbon emission allowance, climate change, crude oil, and 16 agricultural futures.
The contribution of our study is as follows. First, we apply a novel $R^{2}$ decomposed connectedness approach proposed by Balli et al. (2023), which could distinguish the contemporaneous and lagged spillovers. Studying contemporaneous and lagged spillover effects in financial markets is crucial for investment decisions, risk management, and ensuring market stability and healthy development. Second, we conduct detailed analysis on how different agricultural products respond to climate change, carbon emission allowances, and crude oil fluctuations, thereby offering targeted risk management strategies. Third, this paper expands the study on factors influencing commodity prices. It reveals dynamic correlations among crude oil, carbon emission allowance, climate change, and agricultural markets.
The remainder of this paper is organized as follows. Section 2 outlines the methodological framework. Section 3 describes the data. Section 4 presents the empirical results and robustness checks. Section 5 concludes the study.
2 Methodology
Building upon a novel approach for assessing $R^{2}$ decomposed connectedness (Balli et al., 2023; Naeem et al., 2024), this study identifies both contemporaneous and lagged idiosyncratic spillover effects within the crude oil, carbon emission allowance, climate change, and agricultural futures markets. This addresses the earlier research that primarily focused on contemporaneous spillovers, marking a significant expansion of the analysis. As detailed by Balli et al. (2023), the $R^{2}$ decomposed connectedness framework utilizes a VAR($p$) model to incorporate contemporaneous effects:
$\bm{y}_{t}=\sum_{i=0}^{p}\bm{B}_{i}\bm{y}_{t-i}+\bm{u}_{t},\quad\bm{u}_{t}\sim N(\bm{0},\bm{\Sigma})$ (1)
where $\bm{y}_{t}$ and $\bm{u}_{t}$ denote demeaned $K\times 1$ dimensional vectors over time, with $\bm{B}$ and $\bm{\Sigma}$ being dimensional matrices such that $diag(\bm{B}_{0})=0$. Notably, when $p$ is set to zero, the model aligns with the contemporaneous $R^{2}$ decomposed connectedness framework as introduced by Naeem et al. (2024). In other words, Eq. (1) can be delineated as $\bm{y}_{k,t}=\bm{b}_{k}\bm{x}_{k}+\bm{u}_{k,t}$, where $\bm{x}_{t}=[\bm{y}_{t},\bm{y}_{t-1},\ldots,\bm{y}_{t-i},\ldots,\bm{y}_{t-p}]$ constitutes an $K(p+1)\times 1$ dimensional vector, and $\bm{b}_{k}$ is structured as an $1\times K(p+1)$ dimensional vector with its $k$th component being zero.
Typically, the cumulative $R^{2}$ contributions from bivariate linear regressions (BLRs) correspond with the multivariate linear regression (MLR)’s goodness-of-fit only when the variables on the right-hand side (RHS) are not correlated with one another. It is crucial, therefore, to establish a transformation that converts the correlated series $\bm{x}_{k,t}$ [^1] The vector $\bm{x}_{k,t}$ corresponds to $\bm{x}_{t}$but omits the variable on the left-hand side. into a set of orthogonal series. This transformation is facilitated by implementing principal component analysis, where the number of latent factors matches the quantity of RHS variables. As a result, the $R^{2}$ decomposition of an MLR can be formulated as follows:
$\bm{R}_{xx}=\bm{V}\bm{\Lambda}\bm{V}^{\prime}=\bm{C}\bm{C}^{\prime}$ (2)
$\bm{C}=\bm{V}\bm{\Lambda}^{1/2}\bm{V}^{\prime}$ (3)
$\bm{R}^{2,d}=\bm{C}^{2}(\bm{C}^{-1}\bm{R}_{yx})^{2}$ (4)
where $\bm{V}$, $\bm{\Lambda}=diag(\lambda_{1},...,\lambda_{K(p+1)-1})$, alongside $\bm{R}_{xx}$, constitute the eigenvectors, eigenvalues, and Pearson correlation matrices of dimensions $[K(p+1)-1]\times[K(p+1)-1]$ respectively. Additionally, $\bm{R}_{yx}$ and the $\bm{R}^{2,d}$ vectors provide Pearson correlation and $R^{2}$ contribution metrics, each of dimension $[K(p+1)-1]\times 1$. To elaborate, $\bm{R}_{xx}$ captures the Pearson correlation across the RHS variables, whereas $\bm{R}_{yx}$ details the correlation between LHS and RHS variables. The initial set of $K-1$ values in $\bm{R}^{2,d}$ corresponds to the contemporaneous $R^{2}$ contributions, and the subsequent values identify the lagged $R^{2}$ contributions. Thus, the sum of the $\bm{R}^{2,d}$ vector is tantamount to the MLR’s overall goodness-of-fit measure. Finally, stacking the $\bm{R}^{2,d}$ vectors from all $K$ multivariate regressions yields a $K\times K(p+1)$ dimensional $\bm{R}^{2,d}$ decomposition matrix, detailed as $[\bm{R}^{2,d}_{0};...;\bm{R}^{2,d}_{i};...\bm{R}^{2,d}_{p}]$. Consequently, the component $\bm{R}^{2,d}_{0}$ [^2] The diagonal entries of the matrix $\bm{R}^{2,d}_{0}$ are all zero. signifies the contemporaneous spillover effects $\bm{R}^{2,d}_{C}$, whereas the sum $\bm{R}^{2,d}_{1}+\ldots+\bm{R}^{2,d}_{i}+\ldots+\bm{R}^{2,d}_{p}$ represents the aggregate of lagged spillover effects $\bm{R}^{2,d}_{L}$.
Within the Diebold and Yılmaz (2014) framework for measuring connectedness, these components serve as substitutes for the scaled GFEVD matrix, indicating that the total connectedness index (TCI) is the mean $R^{2}$ across all multivariate linear regressions.
$TCI=\frac{1}{K}\sum_{k=1}^{K}R_{k}^{2}=\left(\frac{1}{K}\sum_{k=1}^{K}\sum_{i=1}^{K}\bm{R}_{C,k,i}^{2,d}\right)+\left(\frac{1}{K}\sum_{k=1}^{K}\sum_{i=1}^{K}\bm{R}_{L,k,i}^{2,d}\right)=TCI^{C}+TCI^{L}$ (5)
here, $TCI^{C}$ and $TCI^{L}$ denote the contemporaneous and lagged total connectedness indices, respectively. In addition, since the goodness-of-fit $R_{k}^{2}$ measure ranges from 0 to 1, there is no need for normalization techniques (Chatziantoniou et al., 2021). Further, total directional connectedness TO and FROM others, alongside the net total directional connectedness, are expressed as:
$TO_{i}=\sum_{k=1}^{K}\bm{R}_{C,k,i}^{2,d}+\sum_{k=1}^{K}\bm{R}_{L,k,i}^{2,d}=TO_{i}^{C}+TO_{i}^{L}$ (6)
$FROM_{i}=\sum_{k=1}^{K}\bm{R}_{C,i,k}^{2,d}+\sum_{k=1}^{K}\bm{R}_{L,i,k}^{2,d}=FROM_{i}^{C}+FROM_{i}^{L}$ (7)
$\begin{split}NET_{i}&=TO_{i}-FROM_{i}\\ &=(TO_{i}^{C}+TO_{i}^{L})-(FROM_{i}^{C}+FROM_{i}^{L})\\ &=(TO_{i}^{C}-FROM_{i}^{C})+(TO_{i}^{L}-FROM_{i}^{L})\\ &=NET^{C}_{i}+NET^{L}_{i}.\end{split}$ (8)
Specifically, the $TO_{i}$ index quantifies the proportion of variance in all LHS variables that is attributed to variable $i$. In contrast, the $FROM_{i}$ index measures how much the RHS variables collectively account for the variance in variable $i$. From the perspective of spillover effects, $NET_{i}$ index indicates whether a variable is a net transmitter or receiver of shocks. If series $i$ has a positive (negative) $NET$ value, it is deemed a net transmitter (receiver) of shocks, suggesting it influences other variables more (less) than it is influenced by them. Analogously, the net pairwise directional connectedness ($NPDC_{ij}$), which works in the same manner as the $NET$ connectedness but operates on a mutual, two-way level between pairs of variables. A positive (negative) $NPDC_{ij}$ indicates that variable $j$ accounts for more (less) of the variation in variable $i$ than vice versa.
Table 1: Summary statistics
OilB CarE CliC Soyb Corn Wheat Rice Sugar Cotto Coffe Cocoa Oats Soym Soyo Cattl Milk Lhog Juice Barl
Mean 0.001 0.001 0.000 0.000 0.000 0.000 0.000 0.001 0.000 0.001 0.001 0.000 0.000 0.000 0.001 0.000 0.000 0.002 0.000
Variance 0.001 0.001 0.000 0.000 0.001 0.001 0.000 0.000 0.001 0.000 0.001 0.001 0.005 0.001 0.000 0.000 0.001 0.001 0.000
Skewness -0.730*** -0.577*** -0.020 -1.095*** -1.898*** -0.358*** -1.354*** -0.281*** -3.375*** 0.433*** -1.083*** -2.631*** -0.459*** -0.653*** 0.283*** 3.442*** -2.205*** -0.762*** 0.273***
Ex.Kurtosis 3.370*** 5.491*** 1.779*** 7.565*** 19.347*** 11.338*** 15.520*** 1.780*** 46.721*** 2.946*** 71.817*** 31.460*** 13.158*** 2.701*** 11.247*** 68.055*** 32.473*** 3.566*** 4.315***
JB 433.988*** 1012.642*** 101.843*** 1994.993*** 12503.623*** 4151.900*** 7983.829*** 112.122*** 71680.628*** 303.292*** 166056.293*** 32726.651*** 5596.583*** 289.566*** 4079.262*** 150503.046*** 34545.979*** 483.635*** 608.492***
ERS -10.328*** -8.316*** -9.559*** -12.933*** -12.054*** -11.440*** -12.706*** -5.708*** -13.173*** -7.954*** -1.935* -13.461*** -13.324*** -$1**$1.856*** -8.076*** -11.575 *** -5.442 *** -9.041***
Kendall’s $\rho$ OilB CarE CliC Soyb Corn Wheat Rice Sugar Cotto Coffe Cocoa Oats Soym Soyo Cattl Milk Lhog Juice Barl
OilB 1.000*** 0.065*** 0.099*** 0.172*** 0.125*** 0.119*** 0.089*** 0.136*** 0.179*** 0.041 0.071*** 0.076*** 0.029 0.205*** 0.065*** 0.011 0.041 -0.008 0.014
CarE 0.065*** 1.000*** 0.083*** 0.034 -0.044 -0.035 -0.007 0.047 0.030 0.020 0.028 0.008 0.005 0.073*** 0.034 -0.007 0.013 -0.011 -0.044
CliC 0.099*** 0.083*** 1.000*** 0.039 -0.006 -0.004 -0.003 0.044 0.110*** 0.036 0.051** 0.043 -0.008 0.065*** 0.072*** -0.020 0.028 0.022 -0.036
Soyb 0.172*** 0.034 0.039 1.000*** 0.347*** 0.265*** 0.122*** 0.136*** 0.094*** 0.102*** 0.003 0.161*** 0.379*** 0.351*** 0.055** 0.031 0.036 0.046 0.018
Corn 0.125*** -0.044 -0.006 0.347*** 1.000*** 0.385*** 0.112*** 0.136*** 0.090*** 0.094*** -0.024 0.201*** 0.159*** 0.190*** 0.034 0.012 0.041 0.004 0.036
Wheat 0.119*** -0.035 -0.004 0.265*** 0.385*** 1.000*** 0.131*** 0.148*** 0.125*** 0.139*** 0.011 0.204*** 0.135*** 0.179*** 0.001 0.024 0.048** 0.019 0.035
Rice 0.089*** -0.007 -0.003 0.122*** 0.112*** 0.131*** 1.000*** 0.060** 0.047 0.056** 0.020 0.063** 0.018 0.104*** 0.015 0.019 -0.016 0.008 0.015
Sugar 0.136*** 0.047 0.044 0.136*** 0.136*** 0.148*** 0.060** 1.000*** 0.150*** 0.158*** 0.066*** 0.132*** 0.066*** 0.120*** 0.005 0.030 0.028 0.047 0.027
Cotto 0.179*** 0.030 0.110*** 0.094*** 0.090*** 0.125*** 0.047 0.150*** 1.000*** 0.124*** 0.042 0.109*** 0.071*** 0.104*** 0.057** -0.034 0.059** 0.041 -0.020
Coffe 0.041 0.020 0.036 0.102*** 0.094*** 0.139*** 0.056** 0.158*** 0.124*** 1.000*** 0.089*** 0.107*** 0.060** 0.077*** 0.016 0.004 0.033 0.022 0.036
Cocoa 0.071*** 0.028 0.051** 0.003 -0.024 0.011 0.020 0.066*** 0.042 0.089*** 1.000*** 0.024 -0.001 0.022 0.067*** -0.051** -0.005 -0.047 0.010
Oats 0.076*** 0.008 0.043 0.161*** 0.201*** 0.204*** 0.063** 0.132*** 0.109*** 0.107*** 0.024 1.000*** 0.099*** 0.122*** 0.039 -0.041 0.025 0.011 -0.002
Soym 0.029 0.005 -0.008 0.379*** 0.159*** 0.135*** 0.018 0.066*** 0.071*** $0*$001 0.099*** 1.000*** 0.021 0.021 -0.003 0.004 -0.012 -0.027
Soyo 0.205*** 0.073*** 0.065*** 0.351*** 0.190*** 0.179*** 0.104*** 0.120*** 0.104*** 0.077*** 0.022 0.122*** 0.021 1.000*** 0.066*** 0.015 0.059** 0.043 0.060**
Cattl 0.065*** 0.034 0.072*** 0.055** 0.034 0.001 0.015 0.005 0.057** 0.016 0.067*** 0.039 0.021 0.066*** 1.000*** 0.039 0.064*** 0.006 0.010
Milk 0.011 -0.007 -0.020 0.031 0.012 0.024 0.019 0.030 -0.034 0.004 -0.051** -0.041 -0.003 0.015 0.039 1.000*** 0.016 0.017 -0.007
Lhog 0.041 0.013 0.028 0.036 0.041 0.048** -0.016 0.028 0.059** 0.033 -0.005 0.025 0.004 0.059** 0.064*** 0.016 1.$0**$010 -0.061**
Juice -0.008 -0.011 0.022 0.046 0.004 0.019 0.008 0.047 0.041 0.022 -0.047 0.011 -0.012 0.043 0.006 0.017 -0.010 1.000*** 0.004
Barl 0.014 -0.044 -0.036 0.018 0.036 0.035 0.015 0.027 -0.020 0.036 0.010 -0.002 -0.027 0.060** 0.010 -0.007 -0.061** 0.004 1.000***
Notes: JB and ERS show the results of normality test and unit root test, respectively. * indicates 10% significance level, ** indicates 5% significance level, *** indicates 1% significance level.
3 Data
We select a dataset encompassing various futures markets, including crude oil, carbon emission allowance, climate change, and agriculture, totaling 19 different futures contracts.
The data spans from December 15, 2020, to November 30, 2023, and consists of daily closing prices of continuous contracts, comprising 773 data observation points in total. To ensure data consistency, missing data points are filled with the preceding day’s data. The data is sourced from Wind and Bloomberg.
The starting date of December 15, 2020, was specifically chosen as it marks the launch of the climate change futures.
In the oil category, we included the Brent Oil Future, which is actively traded on the ICE Exchange. For carbon emission allowance, our dataset features the EUA Carbon traded on ICE Futures Europe. In the climate segment, we have incorporated the MSCI World Climate Change Futures, listed on the ICE Exchange, marking an essential inclusion considering the global focus on climate-related financial products.
The agricultural futures in our dataset are quite extensive. They include Soybean, Soybean meal, Soybean oil, Corn, Wheat, Oats, and Rough Rice Futures, all traded on the CBOT Exchange. The ICE Exchange is represented with Sugar #11 Futures, Cotton #2 Futures, Orange juice, Cocoa, and Coffee C Futures; notably, the Coffee C futures contract on ICE serves as the global benchmark for Arabica coffee. Additionally, Live cattle, Milk, Lean hog futures from the CME Exchange are part of our selection. Reflecting the significance of Australia in the global agricultural market, particularly in barley, we also consider Barley Futures listed on the Australian Securities Exchange.
For the sake of simplicity in the presentation of tables and figures, we employ the abbreviations OilB, CarE, CliC, Soyb, Rice, Sugar, Cotto, Coffe, Soym, Soyo, Cattl, Lhog, Juice, and Barl to denote the futures Brent Oil, EUA Carbon, MSCI World Climate Change, Soybean, Rough Rice, Sugar #11, Cotton #2, Coffee, Soybean Meal, Soybean Oil, Live Cattle, Lean Hog, Orange Juice, and Barley, respectively.
Table. 1 reports the descriptive statistic. We find that except coffee, cattle, milk, and barley, other variables are left skewed. Besides that, with the exception of climate change, sugar and soybean oil, others are leptokurtic. The JB and ERS tests indicate that all series are non-normally distributed and significantly stationary. Moreover, the Kendall rank correlation coefficients suggest that most variables are positively correlated.
4 Empirical analysis
4.1 Averaged dynamic connectedness measures
Table 2: Averaged dynamic connectedness.
OilB CarE CliC Soyb Corn Wheat Rice Sugar Cotto Coffe Cocoa Oats Soym Soyo Cattl Milk Lhog Juice Barl FROM
OilB 0.20 2.58 3.13 3.86 2.14 1.89 1.34 3.24 3.44 1.07 1.03 1.25 1.75 4.78 0.81 1.27 1.02 1.20 0.97 36.76
( 0.00, 0.20) ( 1.34, 1.23) ( 2.59, 0.54) ( 3.52, 0.34) ( 1.68, 0.47) ( 1.17, 0.72) ( 1.16, 0.18) ( 2.58, 0.66) ( 3.29, 0.15) ( 0.75, 0.32) ( 0.68, 0.35) ( 0.60, 0.65) ( 0.68, 1.07) ( 4.35, 0.42) ( 0.40, 0.42) ( 0.81, 0.46) ( 0.29, 0.73) ( 0.18, 1.02) ( 0.63, 0.34) (26.70, 10.06)
CarE 1.91 1.23 2.12 0.83 1.62 1.87 1.28 1.75 0.80 0.93 1.23 1.02 0.45 1.16 0.86 0.57 0.57 1.12 1.07 21.15
( 1.45, 0.46) ( 0.00, 1.23) ( 1.49, 0.63) ( 0.56, 0.28) ( 1.23, 0.39) ( 1.57, 0.30) ( 0.49, 0.78) ( 1.44, 0.31) ( 0.47, 0.33) ( 0.20, 0.73) ( 0.46, 0.77) ( 0.38, 0.63) ( 0.17, 0.28) ( 0.93, 0.23) ( 0.66, 0.19) ( 0.26, 0.31) ( 0.43, 0.14) ( 0.72, 0.40) ( 0.89, 0.18) (13.80, 7.35)
CliC 3.10 1.84 0.47 0.74 0.80 0.53 0.57 1.62 2.35 1.05 0.58 0.78 0.64 1.77 1.55 1.15 0.97 1.11 0.92 22.07
( 2.77, 0.33) ( 1.51, 0.33) ( 0.00, 0.47) ( 0.43, 0.31) ( 0.33, 0.47) ( 0.30, 0.23) ( 0.19, 0.38) ( 0.59, 1.02) ( 2.12, 0.24) ( 0.57, 0.49) ( 0.38, 0.20) ( 0.55, 0.23) ( 0.47, 0.17) ( 0.68, 1.08) ( 0.89, 0.67) ( 0.59, 0.56) ( 0.43, 0.54) ( 0.54, 0.57) ( 0.42, 0.50) (13.76, 8.31)
Soyb 3.51 0.52 0.60 0.38 8.91 4.13 1.89 1.51 1.19 1.08 0.73 3.80 5.29 16.29 1.21 0.68 3.38 1.26 0.73 56.71
( 3.33, 0.18) ( 0.42, 0.10) ( 0.33, 0.27) ( 0.00, 0.38) ( 8.35, 0.56) ( 3.60, 0.53) ( 1.54, 0.35) ( 0.93, 0.58) ( 0.51, 0.68) ( 0.74, 0.34) ( 0.56, 0.17) ( 1.58, 2.22) ( 3.65, 1.64) (15.04, 1.25) ( 0.46, 0.75) ( 0.11, 0.57) ( 2.70, 0.67) ( 1.05, 0.22) ( 0.48, 0.25) (45.39, 11.32)
Corn 2.26 1.19 0.49 9.46 1.98 9.42 2.70 1.06 1.03 0.60 0.61 2.67 6.71 3.31 0.75 0.37 0.69 1.28 1.54 46.14
( 1.66, 0.60) ( 1.07, 0.12) ( 0.25, 0.23) ( 8.91, 0.55) ( 0.00, 1.98) ( 8.71, 0.70) ( 2.42, 0.28) ( 0.56, 0.50) ( 0.85, 0.18) ( 0.41, 0.20) ( 0.42, 0.19) ( 2.31, 0.36) ( 5.74, 0.97) ( 2.88, 0.43) ( 0.24, 0.51) ( 0.26, 0.12) ( 0.43, 0.26) ( 0.88, 0.41) ( 1.38, 0.17) (39.37, 6.77)
Wheat 1.97 1.71 0.61 4.30 10.01 0.64 3.74 3.55 1.11 1.44 1.66 3.96 0.92 2.50 1.07 0.80 1.21 0.44 1.13 42.12
( 1.16, 0.81) ( 1.39, 0.32) ( 0.25, 0.36) ( 3.79, 0.51) ( 8.97, 1.04) ( 0.00, 0.64) ( 3.09, 0.65) ( 3.19, 0.36) ( 0.64, 0.47) ( 0.90, 0.54) ( 0.32, 1.35) ( 3.72, 0.23) ( 0.68, 0.24) ( 1.97, 0.53) ( 0.47, 0.60) ( 0.29, 0.51) ( 0.74, 0.46) ( 0.29, 0.16) ( 0.74, 0.38) (32.60, 9.52)
Rice 2.05 0.59 0.44 2.56 3.71 3.61 2.61 1.75 0.93 0.88 1.06 1.98 2.08 1.78 1.43 0.67 0.47 0.55 0.57 27.11
( 1.21, 0.85) ( 0.47, 0.12) ( 0.18, 0.26) ( 1.85, 0.71) ( 2.68, 1.04) ( 3.32, 0.29) ( 0.00, 2.61) ( 1.56, 0.18) ( 0.75, 0.19) ( 0.56, 0.32) ( 0.78, 0.28) ( 1.54, 0.44) ( 0.65, 1.43) ( 0.86, 0.91) ( 0.14, 1.29) ( 0.28, 0.39) ( 0.20, 0.28) ( 0.22, 0.33) ( 0.20, 0.37) (17.43, 9.68)
Sugar 2.86 1.67 1.19 1.92 0.77 3.84 1.91 0.35 1.75 3.97 1.48 3.16 0.54 1.33 1.70 1.10 0.55 1.05 0.54 31.34
( 2.60, 0.26) ( 1.36, 0.32) ( 0.56, 0.64) ( 0.94, 0.99) ( 0.56, 0.22) ( 3.36, 0.48) ( 1.54, 0.37) ( 0.00, 0.35) ( 1.35, 0.40) ( 3.68, 0.29) ( 0.95, 0.53) ( 2.77, 0.39) ( 0.27, 0.27) ( 1.02, 0.31) ( 1.16, 0.54) ( 0.31, 0.79) ( 0.18, 0.37) ( 0.52, 0.53) ( 0.43, 0.11) (23.53, 7.81)
Cotto 3.74 0.60 2.36 0.88 1.16 0.94 1.28 1.77 0.67 2.06 0.89 2.03 0.47 1.00 1.06 1.39 0.77 1.08 1.01 24.50
( 3.45, 0.29) ( 0.46, 0.14) ( 2.10, 0.25) ( 0.53, 0.35) ( 0.94, 0.22) ( 0.67, 0.27) ( 0.76, 0.53) ( 1.39, 0.38) ( 0.00, 0.67) ( 1.48, 0.58) ( 0.39, 0.50) ( 1.85, 0.18) ( 0.23, 0.24) ( 0.74, 0.26) ( 0.36, 0.70) ( 1.08, 0.31) ( 0.62, 0.16) ( 0.79, 0.29) ( 0.28, 0.73) (18.12, 6.38)
Coffe 2.23 0.43 0.93 1.26 0.82 1.31 0.79 4.34 1.83 0.33 1.75 1.95 0.44 0.67 0.72 0.73 1.22 0.94 0.54 22.92
( 0.80, 1.43) ( 0.19, 0.24) ( 0.54, 0.39) ( 0.89, 0.37) ( 0.47, 0.35) ( 0.97, 0.34) ( 0.57, 0.22) ( 3.80, 0.54) ( 1.45, 0.38) ( 0.00, 0.33) ( 1.38, 0.37) ( 1.38, 0.57) ( 0.15, 0.29) ( 0.33, 0.34) ( 0.57, 0.15) ( 0.50, 0.24) ( 0.56, 0.66) ( 0.71, 0.23) ( 0.41, 0.13) (15.68, 7.24)
Cocoa 1.19 1.36 1.39 1.15 1.57 1.12 1.57 1.42 0.80 1.83 1.09 0.81 0.83 1.78 1.18 0.28 1.46 0.84 0.71 21.30
( 0.74, 0.44) ( 0.46, 0.90) ( 0.38, 1.01) ( 0.71, 0.44) ( 0.56, 1.01) ( 0.38, 0.75) ( 0.82, 0.75) ( 1.01, 0.42) ( 0.41, 0.39) ( 1.39, 0.44) ( 0.00, 1.09) ( 0.22, 0.59) ( 0.24, 0.60) ( 0.94, 0.84) ( 0.40, 0.78) ( 0.11, 0.17) ( 0.48, 0.98) ( 0.69, 0.14) ( 0.34, 0.37) (10.28, 11.03)
Oats 1.06 0.68 1.59 2.06 3.50 4.37 1.98 3.01 2.37 1.75 0.31 0.67 0.92 4.72 0.47 0.86 0.48 1.60 1.01 32.75
( 0.64, 0.42) ( 0.34, 0.34) ( 0.53, 1.07) ( 1.87, 0.19) ( 2.51, 0.99) ( 3.93, 0.44) ( 1.49, 0.49) ( 2.74, 0.27) ( 1.75, 0.61) ( 1.29, 0.46) ( 0.21, 0.10) ( 0.00, 0.67) ( 0.51, 0.41) ( 4.36, 0.36) ( 0.35, 0.13) ( 0.36, 0.50) ( 0.28, 0.20) ( 0.57, 1.03) ( 0.74, 0.27) (24.46, 8.28)
Soym 0.92 0.36 0.60 5.05 6.92 1.10 0.82 0.48 0.55 0.48 0.45 0.75 17.14 1.11 0.39 0.19 0.81 0.49 0.77 22.23
( 0.72, 0.20) ( 0.15, 0.21) ( 0.43, 0.18) ( 4.52, 0.53) ( 6.17, 0.75) ( 0.76, 0.34) ( 0.65, 0.17) ( 0.28, 0.20) ( 0.24, 0.31) ( 0.16, 0.33) ( 0.22, 0.23) ( 0.50, 0.25) ( 0.00, 17.14) ( 0.89, 0.22) ( 0.17, 0.22) ( 0.09, 0.10) ( 0.64, 0.17) ( 0.25, 0.24) ( 0.33, 0.44) (17.16, 5.07)
Soyo 4.48 1.03 1.10 16.88 3.37 2.62 1.62 1.47 1.03 0.47 1.26 4.66 1.13 0.34 1.66 0.94 0.63 1.13 0.89 46.38
( 4.22, 0.26) ( 0.78, 0.25) ( 0.60, 0.50) (16.58, 0.30) ( 2.97, 0.40) ( 2.09, 0.54) ( 0.83, 0.79) ( 1.03, 0.44) ( 0.70, 0.33) ( 0.28, 0.19) ( 0.81, 0.45) ( 4.03, 0.63) ( 0.68, 0.45) ( 0.00, 0.34) ( 1.10, 0.56) ( 0.20, 0.74) ( 0.36, 0.27) ( 0.49, 0.64) ( 0.46, 0.43) (38.20, 8.18)
Cattl 0.70 1.23 1.99 0.99 0.88 1.74 0.38 1.54 1.10 0.89 0.82 0.67 0.87 1.85 1.16 1.25 0.87 0.99 0.64 19.39
( 0.43, 0.27) ( 0.66, 0.57) ( 0.91, 1.08) ( 0.52, 0.47) ( 0.31, 0.56) ( 0.57, 1.17) ( 0.16, 0.22) ( 1.22, 0.32) ( 0.37, 0.73) ( 0.58, 0.30) ( 0.41, 0.42) ( 0.38, 0.29) ( 0.21, 0.66) ( 1.22, 0.64) ( 0.00, 1.16) ( 0.85, 0.40) ( 0.64, 0.23) ( 0.32, 0.67) ( 0.31, 0.33) (10.07, 9.32)
Milk 1.76 0.84 2.16 0.40 1.00 0.76 0.63 0.64 1.44 1.06 0.56 0.84 0.36 0.99 2.01 0.20 0.26 1.44 0.30 17.46
( 0.95, 0.81) ( 0.27, 0.57) ( 0.62, 1.55) ( 0.17, 0.23) ( 0.36, 0.64) ( 0.35, 0.41) ( 0.31, 0.32) ( 0.36, 0.28) ( 1.17, 0.27) ( 0.52, 0.53) ( 0.12, 0.45) ( 0.42, 0.42) ( 0.12, 0.25) ( 0.26, 0.74) ( 0.84, 1.17) ( 0.00, 0.20) ( 0.17, 0.09) ( 0.49, 0.95) ( 0.18, 0.13) ( 7.65, 9.81)
Lhog 0.78 0.66 0.83 4.14 1.58 1.11 0.66 0.47 1.03 1.20 1.04 1.51 2.29 0.88 0.93 0.27 0.27 0.76 1.81 21.96
( 0.33, 0.45) ( 0.43, 0.24) ( 0.44, 0.39) ( 3.42, 0.73) ( 0.56, 1.02) ( 0.85, 0.26) ( 0.22, 0.44) ( 0.21, 0.27) ( 0.62, 0.40) ( 0.58, 0.62) ( 0.49, 0.55) ( 0.32, 1.19) ( 0.66, 1.63) ( 0.48, 0.40) ( 0.63, 0.29) ( 0.16, 0.10) ( 0.00, 0.27) ( 0.46, 0.31) ( 1.01, 0.80) (11.85, 10.11)
Juice 0.80 0.95 0.89 1.85 1.72 1.32 1.32 0.84 1.32 1.18 1.21 0.88 0.65 1.23 1.29 1.34 1.75 1.15 1.28 21.82
( 0.20, 0.61) ( 0.73, 0.22) ( 0.53, 0.36) ( 1.56, 0.29) ( 1.04, 0.68) ( 0.36, 0.96) ( 0.24, 1.08) ( 0.55, 0.29) ( 0.80, 0.51) ( 0.71, 0.46) ( 0.69, 0.53) ( 0.62, 0.26) ( 0.29, 0.36) ( 0.68, 0.55) ( 0.32, 0.97) ( 0.47, 0.88) ( 0.46, 1.29) ( 0.00, 1.15) ( 1.02, 0.26) (11.26, 10.57)
Barl 1.22 1.61 1.21 1.23 2.65 3.95 0.70 1.00 1.10 0.78 0.94 1.70 1.47 1.00 0.50 0.31 1.29 1.50 1.13 24.17
( 0.68, 0.54) ( 0.87, 0.74) ( 0.43, 0.78) ( 0.59, 0.64) ( 1.52, 1.14) ( 0.84, 3.11) ( 0.21, 0.48) ( 0.47, 0.53) ( 0.29, 0.81) ( 0.41, 0.37) ( 0.33, 0.62) ( 0.84, 0.86) ( 0.35, 1.12) ( 0.53, 0.47) ( 0.30, 0.20) ( 0.17, 0.15) ( 1.00, 0.29) ( 1.02, 0.49) ( 0.00, 1.13) (10.84, 13.34)
TO 36.54 19.85 23.64 59.57 53.15 45.63 25.17 31.46 25.17 22.72 17.63 34.41 27.82 48.15 19.59 14.18 18.37 18.81 16.44 558.29
( 27.33, 9.20) ( 12.91, 6.94) ( 13.15, 10.49) ( 51.35, 8.21) ( 41.20, 11.95) ( 33.78, 11.85) ( 16.68, 8.49) ( 23.91, 7.55) ( 17.78, 7.38) ( 15.21, 7.51) ( 9.56, 8.06) ( 24.03, 10.38) ( 15.75, 12.07) ( 38.16, 9.99) ( 9.45, 10.14) ( 6.89, 7.29) ( 10.58, 7.79) ( 10.17, 8.63) ( 10.24, 6.20) (388.16, 170.14)
Inc.Own 36.74 21.07 24.11 59.95 55.13 46.27 27.78 31.81 25.83 23.05 18.71 35.08 44.96 48.49 20.74 14.39 18.64 19.96 17.57 TCI
(27.33, 9.41) (12.91, 8.17) (13.15, 10.96) (51.35, 8.59) (41.20, 13.93) (33.78, 12.48) (16.68, 11.10) (23.91, 7.90) (17.78, 8.05) (15.21, 7.84) ( 9.56, 9.15) (24.03, 11.05) (15.75, 29.21) (38.16, 10.33) ( 9.45, 11.29) ( 6.89, 7.49) (10.58, 8.06) (10.17, 9.79) (10.24, 7.34) ($\rm TCI^{C}$, $\rm TCI^{L}$)
NET -0.22 -1.30 1.57 2.85 7.01 3.51 -1.93 0.12 0.67 -0.20 -3.68 1.66 5.59 1.77 0.20 -3.27 -3.59 -3.01 -7.73 29.38
( 0.63, -0.86) (-0.89, -0.41) (-0.60, 2.18) ( 5.96, -3.11) ( 1.82, 5.18) ( 1.18, 2.33) (-0.74, -1.19) ( 0.38, -0.25) (-0.33, 1.00) (-0.47, 0.27) (-0.71, -2.96) (-0.43, 2.10) (-1.41, 7.00) (-0.04, 1.81) (-0.62, 0.82) (-0.76, -2.52) (-1.27, -2.32) (-1.08, -1.93) (-0.60, -7.13) (20.43, 8.95)
  • Notes: $R^{2}$ decomposed measures are based on a 200-day rolling-window VAR model with a lag length of order one (BIC). Values in parentheses represent contemporaneous and lagged effects, respectively.
We illustrate the averaged connectedness measures in Table 2, while values in parentheses illustrate the contemporaneous and lagged $R^{2}$ decomposed measures, respectively. We observe that on average the TCI is equal to 29.38%, indicating that on average 29.38% of variance of the LHS variables can be explained by the RHS variables. Decomposing this metric into its contemporaneous and lagged components, we find that 20.43% are attributed to contemporaneous effects while only 8.95% are caused by lagged dynamics. Furthermore, almost all the contemporaneous FROM and TO connectedness metrics are significantly greater than the corresponding lagged counterparts. However, cocoa, milk, and barley have lagged FROM measures (11.03%, 9.81%, and 13.34%) exceeding their FROM contemporaneous measures (10.28%, 7.65%, and 10.84%). The lagged TO measures for cattle and milk (10.14% and 7.29%, respectively) are both higher than their contemporaneous values (9.45% and 6.89%, respectively).
Additionally, we observe that corn is the primary net transmitter of shocks (7.01%), followed by soybean meal (5.59%), and wheat (3.51%). Several factors might make corn the main net contributor of shocks. First, the close relationship between corn and commodity markets has been widely reported (Kang et al., 2017). Corn is one of the primary raw materials for biofuels, whose demand is closely linked to oil prices. Secondly, corn, requiring a significant amount of water, is particularly sensitive to climate change. Third, corn plays a crucial role in the global food supply chain. Moreover, agricultural activities can also lead to significant carbon emission allowance. Similarly, both soybean meal and wheat are important agricultural products for many countries, which might make them risk contributors. Conversely, barley is the chief net receiver of shocks (-7.73%), with cocoa (-3.68%) and leah hog (-3.59%) following. Besides that, we observe that climate change is a net transmitter (1.57%), while crude oil (-0.22%) and carbon emission allowance (-1.30%) are net risk receivers. In detail, carbon emission (2.12%), cattle (1.99%), and milk (2.16%) receive the largest risk from climate change, while the largest spillovers to climate change are caused by crude oil (3.1%). Barley receives the largest spillovers from wheat (3.95%), with 0.84% being caused by contemporaneous effects while 3.11% originate from lagged spillovers. More specifically, corn and wheat receive major spillovers from each other. The spillover among crude oil and agricultural markets is consistent with the findings of Du et al. (2011).
4.2 Dynamic total connectedness
Figure 1: Dynamic total connectedness. The black margin visualizes the overall dynamic total connectedness while the dynamic contemporaneous and lagged connectedness are illustrated in red and green, respectively
Fig. 1 presents the dynamic total connectedness. We note high market spillovers since 2021. Subsequently, the dynamic total connectedness continued to decrease, implying that the impact of the COVID-19 pandemic is gradually diminishing. However, early in the outbreak of the Russia-Ukraine war, the dynamic total connectedness increased significantly. Afterward, a subsequent decline is observed, with a peak of 28% in June 2023. Since then, the TCI value has reached a plateau of around 25% throughout the period of investigation. Overall, the lagged TCI remained relatively constant except for its increase during the outbreak of the Russia-Ukraine war, which fluctuated around 10% during the entire sample period.
Moreover, we see that the overall and contemporaneous dynamic total connectedness measures behaved extremely similarly. Hence, the contemporaneous interdependencies are more significant than the lagged interdependencies, as they account for the majority of total risk dynamics. The movement in the overall TCI is mainly driven by the contemporaneous dynamics and not by the lagged dynamics, which is in line with the work of Balli et al. (2023).
4.3 Dynamic FROM, TO and NET directional connectedness
Figure 2: FROM connectedness.
Figures 2 and 3 illustrate the FROM and TO directional connectedness, respectively. We find that the contemporaneous spillovers are more pronounced than the lagged spillovers. In Fig. 2, the FROM connectedness for most variables experienced significant fluctuations during the Russia-Ukraine war, especially the crude oil, climate change, wheat and barley. Additionally, soybean has the highest degree of FROM connectedness, followed by soybean oil and wheat.
Figure 3: TO connectedness.
Similarly, Fig. 3 shows that the TO directional connectedness for crude oil, carbon emission allowance, climate change, wheat, and barley shocked during the Russia-Ukraine war. Besides that, the movement of the TO directional connectedness for carbon emission allowance, coffee, cocoa, cattle, milk, lean hog, juice, and barley remain below 30%. Among them, the TO connectedness for cattle is relatively stable.
In addition, a net positive spillover corresponds to a net transmitter of risks to other markets, while a net negative spillover suggests that a particular market is a net receiver of shocks in the system. Fig. 4 presents that the net contributors of shocks in the systems during the sample period mainly include climate change, corn, wheat, oats, soybean meat, and soybean oil, while cocoa, lean hog, juice, and barley mainly act as risk receivers of shocks from other variables. Barley is the biggest receiver of net spillovers. The remaining markets at different periods act as transmitters and receivers alternatively.
A critical look suggests that crude oil behaved as a receiver then switched roles as a transmitter since November 2022. However, carbon emission allowance shows an almost opposite pattern. Moreover, lagged net spillover of most markets exhibits higher similarity to overall net spillover effects, which is in line with Liu et al. (2024). However, different from other markets, soybean initially displayed a higher level of contemporaneous spillovers, followed by a stronger level of lagged correlation.
Figure 4: Net total directional connectedness.
Figure 5: Net pairwise directional connectedness.
4.4 Net pairwise directional connectedness measures
In this subsection, we discuss how one specific market impacts another on a pairwise basis. Fig. 5 shows the net pairwise risk spillovers of the system. By taking a closer look, we find obvious spillovers between climate change and other variables. In the early stages of our selected sample period, climate risk has a net transmitting effect on the crude oil market, which is consistent with Jin et al. (2023). Cocoa, oats, and milk are the recipients of spillover from climate change. On the other hand, climate change receives spillover from the following sectors: corn and soybean oil.
Moreover, we notice that crude oil is a recipient of shocks from soybean, soybean meat, and a contributor of shocks to rice, cotton, and coffee. However, this result differs from the findings of Tiwari et al. (2022). And the net pairwise directional connectedness between crude oil and sugar switches its sign during the sample period, which corroborates the work of Tiwari et al. (2022).
Additionally, we find that lean hog is constantly dominated by soybean, corn, oats, soybean meal, and barley, which accords with Table 2. And barley receives high levels of spillovers from carbon emission allowance, wheat, and corn. Meanwhile, corn and rice transfer risks to the carbon market. We also observe spillovers from cattle, soybean, and soybean meal to rice.
4.5 Net pairwise network analysis
In Fig. 6, we visually illustrate the net pairwise directional connectedness measures. A transmitter is denoted by the blue dot, while a receiver is denoted by the yellow dot. Carbon emission allowance, barley, and rice exhibit negative net spillover effects. Corn and wheat have notable positive risk spillover effects, which is in line with Guo et al. (2023). Moreover, there is a high level of overall and lagged connectedness in wheat and barley. This strong correlation aligns with Banerjee et al. (2024).
Figure 6: Network connectedness. From left to right, they correspond respectively to overall, contemporaneous, and lagged average network dependency.
4.6 Robustness check
Fig. 7 illustrates the robustness checks based on the dynamic total connectedness. We replace the Pearson correlation coefficients with the Spearman and Kendall correlation coefficients. The dynamic TCIs of the three kinds of correlation coefficients are very close, implying that the empirical results are robust.
Figure 7: Robustness check: Dynamic total connectedness. The black line depicts the overall dynamic total connectedness while the dynamic contemporaneous and lagged connectedness are illustrated in red and green, respectively. The three distinct line types represent three different correlation coefficients: Kendall’s $\tau$, Spearman’s $\rho$, and Pearson’s $r$.
5 Conclusion
Climate change has attracted global attention, with profound impacts on agricultural activities, energy extraction, and transportation. To promote a low-carbon economy, protect the environment, and maintain financial stability, it is essential to understand the risk spillovers between climate change, crude oil, carbon emission allowance, and agricultural markets. In this paper, we utilize a novel $R^{2}$ decomposed connectedness approach to investigate the contemporaneous and lagged connectedness across these markets.
The analysis reveals that the total TCI dynamics have been mainly driven by contemporaneous dynamics, which are more volatile than the lagged dynamics. Moreover, the total connectedness has increased due to the Russia-Ukrainian war. Besides that, there is a degree of spillover from climate change to other markets. In particular, the largest transmitter of carbon emission allowance, cattle, and milk is climate change. And there exist heterogenous spillover effects among agricultural markets. In detail, corn is the primary net transmitter of shocks, while barley mainly acts as a net receiver.
Our results have the following implications. By distinguishing between contemporaneous and lagged spillover effects, investors and risk managers can more accurately assess market risks, identify sources of risk and transmission pathways, and develop more effective risk management strategies. The detailed analysis enables market participants to predict future market changes in advance, adjust their investment strategies in a timely manner, and enhance market responsiveness.
Acknowledgements
This work was supported by the Shanghai Planning Office of Philosophy and Social Science (2022EJB006).